OpenStudy (anonymous):

____Cosb= 1/2 [ sin(a+b) + sin (a-b) a. sinb b. cosa c. cosb d. sina

hartnn (hartnn):

sin(a+b) = sin a cos b + cos a sin b sin(a-b) = sin a cos b - cos a sin b

hartnn (hartnn):

add the above 2 equations.

hartnn (hartnn):

what u get ?

OpenStudy (anonymous):

you get... 0 = 0? 0.o wouldn't the minus and plus signs cancil?

hartnn (hartnn):

how u getting 0 anywhere ?

OpenStudy (anonymous):

wait... its a +b ._. sin(a+b) - sin(a-b) =?

hartnn (hartnn):

sin(a+b) = sin a cos b + cos a sin b + sin(a-b) = sin a cos b - cos a sin b ___________________________________________ sin(a+b) +sin (a-b) = 2 sin a cos b + 0 did u get this ?

OpenStudy (nubeer):

subsitute sin (a+b) with sin a cos b + cos a sin b and sin (a-b) with sin a cos b - cos a sin b

OpenStudy (anonymous):

wait .why is there a 0 ..i get everything else in that hartnn

hartnn (hartnn):

cos a sin b - cos a sin b = 0

OpenStudy (anonymous):

oooh ..ty ...now continue please?

hartnn (hartnn):

sin(a+b) +sin (a-b) = 2 sin a cos b 1/2 [sin(a+b) +sin (a-b)] = sin a cos b and you are done.....

OpenStudy (anonymous):

do its d sina right?

hartnn (hartnn):

yes.

hartnn (hartnn):

any more doubts ?

OpenStudy (anonymous):

NO ..and i actually understood that hartnn !! :D

hartnn (hartnn):

thats good :)